A special shoe manufacturer ABC Co. has costs of production as follows :

Quantity                           0     1      2      3     4       5       6           

Total Variable Cost ($)     0     50    70    90   140   200   360

 

a. ABC Co.’s fixed costs are $100. Calculate average fixed costs, average variable costs, average total costs and marginal costs at each level of production.



1
Expert's answer
2021-05-13T14:25:25-0400

a)

"AFC=\frac{TFC}{Q}"


"AVC=\frac{TVC}{Q}"


"ATC=\frac{TC}{Q}"


"MC=TCn-TC(n-1)"